OpenStudy (anonymous):

Completely factor the expression 5x²y + 30xy² + 45yᶾ

OpenStudy (btaylor):

do the prime factorization of all 3 terms: \[\rightarrow (5 \times x \times x \times y) + (2 \times 3 \times 5 \times x \times y \times y) + (5 \times 3 \times 3 \times y \times y \times y)\] You'll notice that all terms have an x , a y, and a 5 as a factor. Pull that out, and you get: \[5xy(x+6y+9y^2)\]